OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, 6 yellow marbles, and 6 green marbles. Four marbles are randomly drawn from the bag, without replacement. The probability of drawing a red, then blue, then green, and finally another red marble is a/b, where a and b are positive, coprime integers. What is the value of a+b?

OpenStudy (kropot72):

Required probability is\[\frac{3}{19}\times \frac{4}{18}\times \frac{6}{17}\times \frac{2}{16}=you\ can\ calculate\]

OpenStudy (anonymous):

waiting for your reply

OpenStudy (kropot72):

@irudayadhason You can cancel and then multiply to find the probability.

OpenStudy (anonymous):

I have no idea about this. Will you please tell me the steps and the answer.

OpenStudy (anonymous):

@kropot72 Thanks

OpenStudy (anonymous):

the answer is 647